IV. An iron wire 1.0 m long with a square cross-section 2.0 mm on a side is used  

              to support a 100 kg load. Find the change in length. 

    

      


1
Expert's answer
2022-05-30T15:18:12-0400
"\sigma=\varepsilon E"

"\frac{F}{A}=\frac{\Delta l}{l_0}E"

"\Delta l=\frac{Fl_0}{AE}=\frac{mgl_0}{AE}"

"\Delta l=\frac{100*9.8*1.0}{4.0*10^{-6}*200*10^9}\\
=1.23*10^{-3}\:\rm m=1.23\: mm"